An industrial imaging system can be purchased for $95,000. At the end of five years, this unit will have a salvage value of $40,000. Alternative, you can lease the machine for five years, with annual payments of $18,000 paid at the beginning of each of the five years. Assuming an APR of 12%, determine whether it is cheaper to buy or to lease the machine
Answer #1

i = 12%

Purchase option:

NPV = -95000 + 40000/(1 + 12%)5 = - $ 72303

Lease option:

NPV = -18000 - 18000/(1 + 12%) - .... - 18000/1.124 = - $ 72672.3

Therefore, it is cheaper to purchase. (higher NPV/lower cost)